You mix 55.0 mL of 0.10 M HCl and 110.0 mL of 0.050 M NH3. What is the pH of the resulting solution? Provide your solution.


1
Expert's answer
2021-04-16T04:59:36-0400

Molarity of HCl = 0.1

Moles of H+= molarity of HCl × volume of sample

= 0.055L × 0.1. = 5.5 × 10-3

Moles of OH- = molarity of NH3 × Volume of sample

= 0.11 × 0.050

= 5.5 × 10-3

= (5.5 × 10-3) /5.5 × 10-3

= 1

= 1/0.165 = 0

PoH= -log(6.1)

= -0.77

= 14--0.77

PH= 14.77